Fermi National Laboratory


Accelerator Update

Friday July 2

The day shift began with Operations monitoring store 3612, with the Antiproton Source (Pbar) off due to work on a Linac RFstation (KRF7), and with the Recycler stashing antiprotons.

Linac technicians completed their work on KRF7 at 8:30 AM. Pbar resumed stacking.

A Booster expert began a cogging study at 10:56 AM. The study lasted twenty minutes.

A Main Injector expert began a study period around 4 PM.

At 9:41 PM, a Booster RF station (BRF1) tripped off. It reset okay.

At 11 PM, the Pbar beam suffered an emittance problem that took about an hour to resolve.

Saturday July 3

The midnight shift began with Operations monitoring store 3612 and the stash, with Pbar stacking, and with MiniBooNE and the Meson MTest experiments taking beam.

At 12:04 AM, a Linac RF station (LRF1) tripped off. It reset okay.

Operations established beam to the MCenter experiment at 12:05 AM.

A Booster kicker (MKS02) tripped off at 12:16 AM. Operators had trouble resetting the device and called in experts. They arrived around 1:30 AM and discovered a problem with the kicker's PFN. They had the station back on by 2 AM and stacking and beam to MiniBooNE and MCenter resumed.

Operations terminated store 3612 at 3:58 AM.

Operations put the Tevatron (TeV) into shot setup at 4:38 AM and disabled the Switchyard (SY) beam.

MTest experimenters finished using the beam at 4:56 AM.

Operations established store 3614 at 6:22 AM with an initial luminosity of 74.18E30.

Operations reestablished beam to MCenter and MTest at 6:35 AM.

Pbar resumed stacking at 6:47 AM. The delay in stacking was due to a problem with the Accumulator septum (A:ISEP1V).

Operators began stashing antiprotons into the Recycler at 9:30 AM.

Stashing ended and stacking resumed at 10:25 AM.

At 11:12 AM, the Linac suffered a trip of two RF stations (LRF1 & 2). Operator investigated and had the stations back up by 11:23 AM.

At 12:26 PM, the Collider Detector at Fermilab (CDF) had to turn off their detector due to colliding beam problems. Operators made some longitudinal damper adjustments.

At 3:44 PM, a SY septum (S:FSEPV) tripped off for the first of seven times in four hours. It reset okay.

At 3:48 PM, a TeV vertical separator sparked and caused some colliding beam problems.

At 4:30 PM, Operations reported that problems with a Pbar Debuncher RF station (DRF2) had been limiting stacking by about 10%. Operators contacted an expert who came in at 4:45 PM.

A Booster RF station (BRF12) tripped off at 5:12 PM.

Operations established beam to MTest at 7:34 PM.

Operations contacted another Pbar RF expert at 8:22 PM that said he'd come in to look at DRF2.

Operations reported at 9:38 PM that BRF12 had tripped off a bit earlier. It reset okay.

At 11:13 PM, Operations reported that BRF12 tripped off again. It reset okay.

MI studies finished at 11:59 PM.

Sunday July 4

The midnight shift began with Operations monitoring store 3614 and the stash, with Pbar stacking, and with MiniBooNE and MTest taking beam.

A Pbar expert managed to get DRF2 up and running normally by 12:16 AM.

At 1:02 AM, a Linac RF station (LRF4) tripped off. It reset okay.

MTest reported that they were done using the beam at 2:43 AM.

At 3:50 AM, Operations reported that BRF12 ran for about twenty minutes and then tripped off again. It tripped again at 4:05 AM and then again at 5:27 AM.

FSEPV tripped off at 5:45 AM. It reset okay.

At 8 AM, Operations reported that BRF12 was off due to an FBS ground fault.

At 8:27 AM, a Debuncher vacuum pump tripped off. Operators had the pump back on twenty-minutes later.

BRF12 tripped off at 9:33 AM and then a few more times throughout the day and early evening.

Operations terminated store 3614 at 9:39 AM.

Operations made the CDF and D-Zero (D0) collision halls ready for access at 9:52 AM. Both experiments were out of access by 11:24 AM.

Operations put the TeV through a dry squeeze at 11:50 AM.

Operations put the TeV into shot setup and then established store 3616 at 2:55 PM with an initial luminosity of 75.16E30. Pbar resumed stacking.

Operations began stashing antiprotons to the Recycler at 5:09 PM. Pbar resumed stacking at 5:50 PM.

A Pbar expert began investigating the lower that expected antiproton production at 7:39 PM.

At 11:08 PM, a lightning strike caused a glitch that tripped off the MI beam safety system permit. LRF1 & 2 both tripped off. Both stations reset okay.

Monday July 5

The midnight shift began with Operations monitoring store 3616 and the stash, with Pbar stacking, and with MiniBooNE and Meson taking beam.

At 2:24 AM, the TeV suffered a sector D0 quench. The D0 service building fire alarm also went off. In addition, the power tripped off for this building. This was a multi-house quench. After the Fire Department said everything was okay, the Duty Electrician began investigating the problem.

BRF13 tripped off at 2:35 AM. It reset okay.

At 3:15 AM, the E0 service building had a fire alarm due to a cryo system cold compressor that smoked. Operators reported that nine compressors tripped off around the TeV ring during the quench. CHL reported at 4 AM that they were having trouble getting some of the compressor back on.

At 5:46 AM, a LCW water pump for LRF5 tripped off. Operators had to add water to the cavity system. They had the pump and the station back on by 6:35 AM. Pbar resumed stacking.

Operations reestablished beam to MTest at 6:43 AM.

LRF5 tripped off at 7:34 AM. Operators worked at getting the station working again.

BRF16 tripped off at 8 AM. It reset okay.

Plans for Monday Morning
The plan for today is to fix the LRF5 station and then to continue with stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/5/2004   email Fermilab